# Common and Rare Fundus Diseases Identification Using Vision-Language Foundation Model with Knowledge of Over 400 Diseases

### v1

@@ Abstract
The current retinal artificial intelligence models were trained using data with a limited category of diseases and limited knowledge. In this paper, we present a retinal vision-language foundation model (RetiZero) with knowledge of over 400 fundus diseases. Specifically, we collected 341,896 fundus images paired with text descriptions from 29 publicly available datasets, 180 ophthalmic books, and online resources, encompassing over 400 fundus diseases across multiple countries and ethnicities. RetiZero achieved outstanding performance across various downstream tasks, including zero-shot retinal disease recognition, image-to-image retrieval, internal domain and cross-domain retinal disease classification, and few-shot fine-tuning. Specially, in the zero-shot scenario, RetiZero achieved a Top5 score of 0.8430 and 0.7561 on 15 and 52 fundus diseases respectively. In the image-retrieval task, RetiZero achieved a Top5 score of 0.9500 and 0.8860 on 15 and 52 retinal diseases respectively. Furthermore, clinical evaluations by ophthalmology experts from different countries demonstrate that RetiZero can achieve performance comparable to experienced ophthalmologists using zero-shot and image retrieval methods without requiring model retraining. These capabilities of retinal disease identification strengthen our RetiZero foundation model in clinical implementation.
@@ 1 Introduction
Blindness and visual impairment represent a significant disease burden globally, impacting millions of individuals. Detection and timely treatment of ocular conditions, such as retinal and optic nerve diseases, are crucial for reducing severe and permanent damage. However, the insufficient availability of ophthalmic medical resources severely limits the prompt screening and treatment of retinal diseases. In recent years, artificial intelligence (AI)-based retinal disease screening systems have alleviated the workload on healthcare professionals to some extent, providing crucial technological support for the timely screening of retinal disease patients and referring for treatment. Nevertheless, most previous AI-based methods were customized for specific diseases, such as diabetic retinopathy (DR) [1, 2], glaucoma [3, 4], and retinopathy of prematurity (ROP) [5, 6]. Although several methods have been proposed for simultaneously screening multiple retinal diseases with promising performance [7, 8], most current AI models for ocular disease screening were trained on task-specific datasets, leading to inevitable errors in prediction when new data (ex., images acquired by different camera) or changes in tasks (ex., introducing new or rare categories). Furthermore, due to limited healthcare resources and the varying prevalence of retinal disease, collecting comprehensive datasets covering all retinal abnormalities is time-consuming and challenging. Consequently, most AI models were trained on limited data and disease categories, restricting their feature representation. Applying these models to different real-world settings or tasks requires extensive retraining with large datasets. Moreover, data quality and labeling issues further limit the widespread adoption of AI models in ophthalmic clinical settings.
@@
Driven by abundant of big data and robust computing hardware, large foundation models (LFMs) have excelled in computer vision tasks [9, 10]. Pre-trained on massive datasets, LFMs provide rich feature support for downstream tasks like object detection [11], few-shot recognition [12], and zero-shot [13], etc. The first ophthalmic LFM, RETFound [14], introduced in 2023, was trained on large, unannotated retinal images using masked autoencoder (MAE) framework [15]. It provides rich feature support and improves the performance of downstream tasks, including internal domain and cross-domain retinal disease classification, few-shot learning, and prediction of systemic diseases. However, such an approach can hinder the model’s capacity to align feature information with labels in downstream tasks. In contrast, the Foundation LAnguage-Image model of the Retina (FLAIR), a Contrastive Language-Image Pre-training (CLIP) -based LFMs enhance feature representation by aligning text descriptions with image features, improving feature-label alignment but struggling with complex semantic features in medical imaging [10]. Current LFMs for ophthalmic imaging are pre-trained on extensive yet categorically limited datasets. Therefore, developing LFMs with comprehensive ophthalmic disease knowledge is crucial for representing complex retinal features and enhancing downstream task performance. However, collecting massive and diverse ophthalmic data that covers a wide range of retinal diseases for pretraining remains a significant challenge.
@@
To address these problems and challenges, we collected 341,896 fundus images paired with text descriptions from 29 publicly available datasets (containing 303,124 fundus images with labels), 180 ophthalmic books (2,3328 fundus images with text descriptions), and online resources (15,544 fundus image with text descriptions), encompassing over 400 retinal and optic nerve diseases across multiple countries/regions and ethnicities. As shown in Figure  1, RetiZero is based on a contrastive vision-language pretraining framework that integrates MAE-based pretraining knowledge and low-rank training methods. Moreover, we introduced an uncertainty vision-language feature calibration method using Dirichlet reparameterization within the contrastive vision-language pretraining framework, to further better align vision-language features in the high-dimensional embedding space. Consequently, RetiZero achieved superior performance in various downstream tasks, including zero-shot fundus disease recognition, image-to-image fundus disease retrieval, internal domain retinal disease identification, few-shot fine-tuning, and cross-domain fundus disease identification.
@@ 2.1 Zero-shot fundus disease recognition
The biggest advantage of RetiZero is the capability of zero-shot learning, which enables RetiZero to recognize fundus diseases using only textual prompts, without needing to retrain or fine-tune the model with labeled fundus images. As shown in Figure 2 (a), RetiZero achieved overall Top 1, Top 3, and Top 5 scores of 0.4421, 0.7024, 0.8404, respectively, for recognizing 15 common fundus diseases and normal condition of 30,089 fundus images (Eye-15 dataset), improved by 25.52%, 15.68%, and 15.61% over FLAIR (a recent VLM foundation model) [16], respectively. Furthermore, in the analysis of individual disease, RetiZero shows outstanding zero-shot capability in identifying most categories, especially for glaucoma (Top1: 0.7477, Top3: 0.9292, and Top5: 0.9718), Retinal Detachment (Top1: 0.7624, Top3: 0.9060, and Top5: 0.9628), and Retinitis Pigmentosa (Top1: 0.8190, Top3: 0.9533, and Top5: 0.9737) (Supplementary Figure 1). Additionally, to further validate RetiZero’s zero-shot capability in more challenging clinical scenarios, we collected a more demanding dataset named EYE-52. This dataset comprises 7,007 fundus images from various ophthalmology clinics, covering 52 fundus diseases, many of which are extremely rare in clinical practice. The incidence/prevalence of each category in the EYE-52 dataset was shown in Supplementary Table 3. As depicted in Figure 2 (b), RetiZero achieved overall Top1, Top3, and Top5 scores of 0.3595, 0.6259, and 0.7561, respectively, for recognizing the 52 types of fundus diseases in a zero-shot manner, providing superior performance compared FLAIR [16] (Top1: 0.0915, Top3: 0.2626, and Top5: 0.3398) and Random recognizing (Top1: 0.0294, Top3: 0.0882, and Top5: 0.1471). Furthermore,
@@
Figure 1: Overview of the framework. Datasets for RetiZero pretraining: The RetiZero model was pre-trained using data from three primary sources: public datasets, ophthalmologic books, and online resources. We assembled a team of 10 ophthalmologists for manual data collection, with an additional two members dedicated to data cleaning tasks. This involved downloading images and corresponding labels from public datasets, extracting images and text descriptions from ophthalmic books, and downloading retinal diseases-relevant images and text descriptions from online resources. RetiZero combines the strengths of self-supervised learning based on the MAE architecture and contrastive learning from the CLIP architecture. Moreover, we introduce an uncertainty vision-language feature calibration method into the contrastive vision-language pretraining framework, to further calibrate visual-language features in the high-dimensional embedding space. Task I: Zero-shot fundus disease recognition. Task II: Fundus disease identification by image-to-image retrieval. Task III: Internal domain retinal disease identification. "Internal domain" means that we fine-tuned and tested the model using the data with similar feature distribution. Task IV: Few-shot fine-tuning. We fine-tuned the model using a few training samples to evaluate RetiZero’s performance in identifying fundus diseases with very limited training data. Task V: Cross-domain fundus disease identification. "Cross-domain" means that we fine-tuned and tested the model using the data with different feature distributions.
@@
RetiZero demonstrated superior zero-shot performance, especially for recognizing some rare fundus diseases in clinical practice. For instance, RetiZero achieved Top 1, Top 3, and Top 5 scores of 0.6163, 0.7907, and 0.8605, respectively, for identifying Bietti Crystalline dystrophy. For the recognition of chorioretinal coloboma, the Top 1, Top 3, and Top 5 scores were 0.5085, 0.8079, and 0.9153, respectively (Supplementary Figure 2). Figure 2 (c) shows the Top5 prediction results provided by RetiZero and FLAIR for three rare disease samples. It can be seen that RetiZero’s Top5 predictions include the correct disease, further demonstrating the outstanding performance of RetiZero in screening rare diseases. More details on the rest of the 52 disease categories can be found in Supplementary Figure 2.
@@ 2.3 Clinical evaluation by ophthalmology experts from different countries
To comprehensively evaluate RetiZero’s capability in fundus disease recognition without retraining the model, we randomly selected two samples from each category of the EYE-52 dataset, creating a new subset called EYE52-sub with a total of 104 instances. 14 ophthalmic experts from Singapore, the United States, and China were invited to diagnose the 104 samples in the EYE52-sub dataset. Among them, four ophthalmic experts have 3 to 5 years of clinical experience, six have 5 to 10 years of clinical experience, and four have more than 10 years. Specifically, we developed an online fundus image reading system and uploaded the 104 samples to the server, as shown in Figure 3. Mimicking the zero-shot setup, we provided 52 disease options on the webpage as prompts. During the image reading process, the clinicians selected diagnostic results from the 52 disease categories based on the image content. Finally, each expert was asked to assess their confidence in their diagnostic results. As shown in Figure 3, the diagnostic accuracy of the 14 ophthalmologists ranges from 0.365 to 0.788, and the median values between ophthalmologists is 0.582, while RetiZero’s zero-shot Top1, Top3, and Top5 accuracies are 0.308, 0.635, and 0.798, respectively. Therefore, RetiZero’s zero-shot Top3 performance is comparable to that of most ophthalmologists, and its Top5 performance surpasses that of all ophthalmologists. Furthermore,the fundus disease identification performance by image-to-image retrieval for RetiZero achieved Top1, Top3, and Top5 accuracies of 0.6837, 0.7449, and 0.7959, respectively. Therefore, RetiZero’s Top1 accuracy in identifying fundus diseases through image-to-image retrieval surpasses that of most ophthalmologists. These experimental results further demonstrate that RetiZero can achieve performance comparable to experienced ophthalmologists through zero-shot and image retrieval methods without retraining the model.
@@ 2.4 Internal domain retinal disease identification
Figure 3: Clinical evaluation. (a) Ophthalmologist diagnostic results, top1, top3 and top5 performance for zero-shot and image-to-image retrieval. (b) Our developed online fundus image reading system.

### v2

@@ Abstract
Previous foundation models for retinal images were pre-trained with limited disease categories and knowledge base. Here we introduce RetiZero, a vision-language foundation model that leverages knowledge from over 400 fundus diseases. For RetiZero’s pre-training, we compiled 341,896 fundus images paired with text descriptions, sourced from public datasets, ophthalmic literature, and online resources, encompassing a diverse range of diseases across multiple ethnicities and countries. RetiZero exhibits superior performance in several downstream tasks, including zero-shot disease recognition, image-to-image retrieval, and internal- and cross-domain disease identification. In zero-shot scenarios, RetiZero achieves Top5 accuracy scores of 0.8430 for 15 fundus diseases and 0.7561 for 52 fundus diseases. For image retrieval, it achieves Top5 scores of 0.9500 and 0.8860 for the same disease sets, respectively. Clinical evaluations show that RetiZero’s Top3 zero-shot performance surpasses the average of 19 ophthalmologists from Singapore, China, and the United States. Furthermore, RetiZero significantly enhances clinicians’ accuracy in diagnosing fundus disease. These findings underscore the value of integrating the RetiZero foundation model into clinical settings, where a variety of fundus diseases are encountered.
@@ 1 Introduction
Blindness and visual impairment represent a significant disease burden globally, impacting millions of individuals across all populations. Detection and timely treatment of ocular conditions, such as retinal and optic nerve diseases, are crucial for reducing severe and permanent damage. However, the insufficient availability of ophthalmic medical resources severely limits the prompt screening and treatment of retinal diseases with vast regional differences in many parts of the world. In recent years, artificial intelligence (AI)-based retinal disease screening systems have been proposed and achieved promising performance in retinal disease detection and referring patients for treatment. Nevertheless, most previous AI-based methods were customized for specific diseases, such as diabetic retinopathy (DR) [1, 2], glaucoma [3, 4], and retinopathy of prematurity (ROP) [5, 6]. Although several methods have been proposed for simultaneously screening multiple retinal diseases with promising performance [7, 8, 9], most current AI models for ocular disease screening were trained on task-specific datasets, leading to inevitable errors in prediction when there were new data (e.g., images acquired by different camera) or changes in tasks (e.g., introducing new or rare categories). Furthermore, due to limited healthcare resources and the varying prevalence of retinal disease, collecting comprehensive datasets covering all kinds of retinal abnormalities is time-consuming and challenging. Consequently, most AI models were trained on limited data and disease categories, restricting their feature representation. Applying these models to different real-world settings or tasks requires extensive retraining with large datasets. Moreover, data quality and labeling issues further limit the widespread adoption of AI models in ophthalmic clinical settings, especially from a global perspective.
@@
Driven by the abundance of big data and robust computing hardware, large foundation models (LFMs) have excelled in computer vision tasks  [10, 11]. Pre-trained on massive datasets, LFMs provide rich feature support for downstream tasks, such as object detection [12], few-shot recognition [13], and zero-shot [14], etc. The first ophthalmic LFM, RETFound [15], introduced in 2023, was trained on large, unannotated retinal images using masked autoencoder (MAE) framework [16]. It provides rich feature support and improves the performance of downstream tasks, including internal domain and cross domain retinal disease classification, few-shot learning and prediction of systemic diseases. However, such an approach can hinder the model’s capacity to align feature information with labels in downstream tasks. In contrast, the Foundation LAnguage-Image model of the Retina (FLAIR), a Contrastive Language-Image Pre-training (CLIP) -based LFMs enhance feature representation by aligning text descriptions with image features, improving feature-label alignment but having difficulties with complex semantic features in medical imaging [11]. Current LFMs for ophthalmic imaging are pre-trained on extensive yet categorically limited datasets. Therefore, developing LFMs with comprehensive ophthalmic disease knowledge would be crucial for representing complex retinal features and enhancing downstream task performance. However, collecting massive and diverse ophthalmic data that covers a wide range of retinal diseases for pretraining remains a significant challenge.
@@
To address these problems and challenges, we collected 341,896 fundus images paired with text descriptions from 29 publicly available datasets (containing 303,124 fundus images with labels), 180 ophthalmic books (2,3328 fundus images with text descriptions), and online resources (15,544 fundus image with text descriptions), encompassing over 400 retinal and optic nerve diseases across multiple countries/regions and ethnicities. As shown in Figure  1, RetiZero is based on a contrastive vision-language pretraining framework that integrates MAE-based pretraining knowledge and low-rank training methods. Moreover, we introduced an uncertainty vision-language feature calibration method using Dirichlet reparameterization within the contrastive vision-language pretraining framework, to further align vision and language features in the high-dimensional embedding space. Consequently, RetiZero achieved superior performance in various downstream tasks, including zero-shot fundus disease recognition, image-to-image fundus disease retrieval, internal domain retinal disease identification, few-shot fine-tuning, and cross-domain fundus disease identification.
@@ 2.1 Zero-shot fundus disease recognition
The biggest advantage of RetiZero is the capability of zero-shot learning, which enables RetiZero to recognize fundus diseases using only textual prompts, without needing to retrain or fine-tune the model with labeled fundus images. As shown in Figure 2 (a), RetiZero achieves overall Top 1, Top 3, and Top 5 scores of 0.4421, 0.7024, 0.8404, respectively, for recognizing 15 common fundus diseases and normal condition of 30,089 fundus images (Eye-15 dataset), improved by 25.52%, 15.68%, and 15.61% over FLAIR (a recent VLM foundation model) [16], respectively. Furthermore, in the analysis of individual diseases, RetiZero shows remarkable zero-shot capability in identifying most categories, especially for glaucoma (Top1: 0.7477, Top3: 0.9292, and Top5: 0.9718), retinal detachment (Top1: 0.7624, Top3: 0.9060, and Top5: 0.9628), and retinitis pigmentosa (Top1: 0.8190, Top3: 0.9533, and Top5: 0.9737) (Supplementary Figure 1). In addition, to further validate RetiZero’s zero-shot capability in more challenging clinical scenarios, we collected a more demanding dataset named EYE-52. This dataset comprises of 7,007 fundus images from various ophthalmology clinics, covering 52 fundus diseases, many of which are extremely rare in clinical practice. The incidence/prevalence of each category in the EYE-52 dataset were shown in Supplementary Table 3. As depicted in Figure 2 (b), RetiZero achieved overall Top 1, Top 3, and Top 5 scores of 0.3595, 0.6259, and 0.7561, respectively, for recognizing these 52 types of fundus diseases in a zero-shot manner, providing superior performance compared FLAIR [17] (Top1: 0.0915, Top3: 0.2626, and Top5: 0.3398) and Random recognizing (Top1: 0.0294, Top3: 0.0882, and Top5: 0.1471). Furthermore,
@@
Figure 1: Overview of the framework. Datasets for RetiZero pretraining: The RetiZero model was pre-trained using data from three primary sources: public datasets, ophthalmology books, and online resources. We assembled a team of 10 ophthalmologists for manual data collection, with an additional two ophthalmologists dedicated to data cleaning tasks. This involved downloading images and corresponding labels from public datasets, extracting images and text descriptions from ophthalmic books, and downloading retinal diseases-relevant images and text descriptions from online resources. RetiZero combines the strengths of self-supervised learning based on the MAE architecture and contrastive learning from the CLIP architecture. Moreover, we introduce an uncertainty vision-language feature calibration method into the contrastive vision-language pretraining framework, to further calibrate visual-language features in the high-dimensional embedding space. Task I: Zero-shot fundus disease recognition. Task II: Fundus disease identification by image-to-image retrieval. Task III: Internal domain retinal disease identification. "Internal domain" means that we fine-tuned and tested the model using the data with similar feature distribution. Task IV: Few-shot fine-tuning. We evaluate RetiZero’s performance in identifying fundus diseases with very limited training data. Task V: Cross-domain fundus disease identification. "Cross-domain" means that we fine-tuned and tested the model using the data with different feature distributions.
@@
RetiZero demonstrated superior zero-shot performance, especially for recognizing some rare fundus diseases in clinical practice. For instance, RetiZero achieved Top 1, Top 3, and Top 5 scores of 0.6163, 0.7907, and 0.8605, respectively, for identifying Bietti Crystalline dystrophy. For the recognition of chorioretinal coloboma, the Top 1, Top 3, and Top 5 scores were 0.5085, 0.8079, and 0.9153, respectively (Supplementary Figure 2). Figure 2 (c) shows the Top 5 prediction results provided by RetiZero and FLAIR for three rare disease samples. Notably, RetiZero’s Top 5 predictions include the correct disease, further demonstrating the excellent performance of RetiZero in screening rare diseases. More details on the rest of the 52 disease categories can be found in Supplementary Figure 2.
@@ 2.3 Clinical evaluation by ophthalmology experts from different countries
To comprehensively evaluate RetiZero’s capability in fundus disease recognition without retraining the model, we randomly selected two samples from each category of the EYE-52 dataset, creating a new subset called EYE52-sub with a total of 104 instances. We invited 19 ophthalmologists from Singapore, the United States, and China, crossing 12 different institutions, to diagnose the 104 samples in the EYE52-sub dataset. Among them, seven ophthalmologists have 3 to 5 years of clinical experience, seven have 5 to 10 years of clinical experience, and five have more than 10 years. Specifically, we developed an online fundus image reading system and uploaded the 104 samples to the server, as shown in Figure 3. Mimicking the zero-shot setup, we provided 52 disease options on the webpage as prompts. During the image reading process, the clinicians selected diagnostic results from the 52 disease categories based on the image content. Finally, each ophthalmologist was asked to assess their confidence in their diagnostic results. As shown in Figure 3, the diagnostic accuracy of the 19 ophthalmologists ranges from 0.337 to 0.788, and the median values between ophthalmologists is 0.582, while RetiZero’s zero-shot Top1, Top3, and Top5 accuracies are 0.308, 0.635, and 0.798, respectively. Therefore, RetiZero’s zero-shot Top3 performance is comparable to that of most of the ophthalmologists, and its Top5 performance surpasses that of all ophthalmologists. Furthermore, the fundus disease identification performance by image-to-image retrieval for RetiZero achieved Top1, Top3, and Top5 accuracies of 0.6837, 0.7449, and 0.7959, respectively. Therefore, RetiZero’s Top1 accuracy in identifying fundus diseases through image-to-image retrieval surpasses that of most ophthalmologists. These experimental results further demonstrate that RetiZero can achieve performance comparable to experienced ophthalmologists through zero-shot and image retrieval methods without retraining the model.
To further verify whether our RetiZero can assist doctors in improving the accuracy of fundus disease diagnosis, we engaged these 19 ophthalmologists in a second round of clinical evaluations, separated by a one-week wash-out period. In this round, identical questions from the first round were presented, but the sequence was randomized, and for each question, RetiZero provided its top five most probable categories. This methodology ensured a comparative analysis of diagnostic performance with and without the assistance of RetiZero’s predictions. As shown in Figure 3(a), the performance of 18 out of the 19 ophthalmologists improved after introducing RetiZero for assisted diagnosis. Their accuracy range of ophthalmologists increased from 0.337 0.788 to 0.529 0.856. This indicates that RetiZero can effectively enhance diagnostic accuracy for ophthalmologists in most cases. Specifically, as shown in Figure 3(b), among these questions, 83 (79.8%) had the correct diagnosis within RetiZero’s top five reference answers, indicating that the majority of RetiZero’s top five suggestions are accurate. Out of 1,976 total responses (104 questions×19 doctors) across both rounds, 1569 responses (79.4%) remained unchanged, while 407 responses (20.6%) were modified in the second round. Of these 407 modified responses, 279 (68.6%) were changed from incorrect to correct, showing that the majority of AI-assisted modifications helped doctors correct previous misdiagnoses. Compared to the first round without RetiZero assistance, the overall accuracy in the second round, which included RetiZero’s assistance, increased by 7.6%. This indicates that RetiZero can help ophthalmologists improve their diagnostic accuracy for fundus diseases, highlighting the significant clinical importance of RetiZero in aiding diagnosis. Statistical analysis of answer modifications between two rounds of clinical evaluations revealed a significant difference between the groups (p<0.001), indicating that RetiZero’s reference answers effectively guide doctors in correcting their diagnoses. The average diagnostic accuracy of the doctors improved markedly from 55.2% to 62.8% (Wilcoxon signed-rank test, p<0.001), reflecting an overall increase of 13.8%. Moreover, as shown in Figure 3(b), we categorized the ophthalmologists by years of experience: Junior ($\leq$5 years, 7 doctors), Senior (5-10 years, 7 doctors), and Expert (>10 years, 5 doctors). Their average first-round accuracies were 48.5%, 56.9%, and 62.3%, respectively, improving to 57.4%, 63.9%, and 69.0% in the second round, with respective increases of 18.4%, 12.3%, and 10.8%. This demonstrates that RetiZero-assisted diagnosis enhances accuracy across all experience levels, with Juniors benefiting the most. Additionally, doctors’ confidence in their diagnoses increased from an average of 2.7 (Low to Moderate) in the first round to 3.0 (Moderate) in the second round, suggesting that RetiZero assistance not only improves accuracy but also boosts diagnostic confidence.
Figure 3: Clinical evaluation. (a) Ophthalmologist diagnostic results, Top1, Top3, and Top5 performance for zero-shot and image-to-image retrieval. (b) Details for clinical evaluation. (c) Online fundus image reading system without RetiZero assistance. (d) Online fundus image reading system with RetiZero assistance.
We further conducted an in-depth analysis of the correlation between the position of the correct diagnosis within RetiZero’s Top 5 predictions and the modifications made by ophthalmologists to their diagnoses. The top five results provided by RetiZero were scored as follows: correct diagnoses appearing in positions 1, 2, 3, 4, and 5 were assigned 5, 4, 3, 2, and 1 points, respectively, while those not appearing in the top five were assigned 0 points. Therefore, the top ranking score of each case can be calculated, which we used to assess the contribution of the correct diagnosis’s ranking within the top five to the diagnostic accuracy. Moreover, we assessed the modifications made by ophthalmologists to their answers. Assuming an ophthalmologist’s answers in the first and second rounds for the same case were denoted as (x, y), the scoring method was as follows: (True, True) = 0, (False, False) = 0, (True, False) = -1, (False, True) = 1. By analyzing the modification behaviors of 19 ophthalmologists, we calculated the response modification score for each case, which reflected the ophthalmologists’ thought process and decision-making during the answering process. Finally, we conducted a Spearman correlation analysis between top-ranking scores from RetiZero and the response modification scores of ophthalmologists. The results showed a correlation coefficient of r=0.614, p<0.001. These results demonstrate that the higher the correct diagnosis is ranked by RetiZero, the higher the diagnostic accuracy of the ophthalmologists for that case.